A retired couple invested part of $11,400 at 6% interest and the rest at 7.5%. If their annual income from these investments is $783. Find the invested part at each rate?
Answer by stanbon(75887)

A retired couple invested part of $11,400 at 6% interest and the rest at 7.5%. If their annual income from these investments is $783. Find the invested part at each rate?
------
Equations:
x + y = 11,400
0.075x + 0.06y = 783
-----------------------------
Multiply thru 1st equation by 75
Multiply thru 2nd equation by 1000
-----
75x + 75y = 75*11400
75x + 60y = 783000
-----
Subtract and solve for "y":
15y = 72000
-----
y = $4800 (amt invested at 6%)
----------------
Solve for "x":
x + y = 11,400
x + 4800 = 11,400
x = $6600 (amt invest at 7.5%)
================================
Cheers,
Stan H.
=============